# Adaptive Prediction Ensemble:

### v1

@@ Abstract
Deep learning-based trajectory prediction models for autonomous driving often struggle with generalization to out-of-distribution (OOD) scenarios, sometimes performing worse than simple rule-based models. To address this limitation, we propose a novel framework, Adaptive Prediction Ensemble (APE), which integrates deep learning and rule-based prediction experts. A learned routing function, trained concurrently with the deep learning model, dynamically selects the most reliable prediction based on the input scenario. Our experiments on large-scale datasets, including Waymo Open Motion Dataset (WOMD) and Argoverse, demonstrate improvement in zero-shot generalization across datasets. We show that our method outperforms individual prediction models and other variants, particularly in long-horizon prediction and scenarios with a high proportion of OOD data. This work highlights the potential of hybrid approaches for robust and generalizable motion prediction in autonomous driving.
@@ I Introduction
Trajectory prediction is critical for safe and reliable autonomous vehicle systems. Existing prediction algorithms [1, 2, 3, 4] have achieved high accuracy on real-world scenarios, such as real traffic datasets. However, most of these algorithms only work best for in-distribution scenarios. Intuitively, traffic scenarios in different cities of the same country should not possess drastic differences, and human driving skills including their prediction and judgment, are not significantly affected. This is unfortunately not the case for deep learning-based prediction algorithms [5, 6, 7]. If they are applied to out-of-distribution (OOD) scenes in a zero-shot manner, such as predicting vehicle trajectories from a different dataset than the training dataset, the performance will drop dramatically even though the input representation and format are the same. In some cases, a deep learning-based prediction algorithm is not even as good as a simple constant velocity model, as shown in Fig. 1. This is unfortunately a largely under-explored topic. One natural way is to combine the prediction from different sources, which resembles the mixture of experts. As far as we know, we are the first to explore concrete methods to improve OOD generalization to different datasets than training.
@@
Figure 1: Illustration of the motivation of improving prediction algorithm by Adaptive Prediction Ensemble. (a) An example scenario where vanilla rule-based prediction algorithm outperforms deep neural network prediction algorithm (MTR [2]). (b) A comparison of the error (minADE) between deep NN and rule-based prediction. The rule-based method outperforms deep NN in a considerable amount of scenarios, which are the ones below the red line.
@@
Mixture of Experts (MoEs) [8] has gained popularity, especially after the great success of Large Language Models. Most of the prior work showed MoEs can reach faster inference [9, 10] compared to dense models with the same number of parameters, and can also be pre-trained faster [11, 12]. While they focused on MoEs’ advantage over a comparable dense model in size, we are more interested in investigating the generalization ability improvements upon a single expert model. We generally find that deep learning prediction models tend to overfit their training dataset, making zero-shot performance unacceptable. Incorporating a fleet of deep learning prediction experts or adopting a similar size large dense model would not solve the problem, since increasing the model capacity would not mitigate the overfitting problem if not making it worse. Therefore, we propose to employ a rule-based prediction expert as an anomaly-handling strategy for deep learning prediction experts, in light of the insight that rule-based prediction could be more reliable in long-tail cases of deep learning prediction experts.
@@
There are other existing methods for domain generalization that usually handle the problem by data manipulation [13], representation learning [14, 15], or specially designed learning strategy [16, 17] or inference workflow [18]. If one aims to improve the generalization upon an existing prediction model with prior methods, it is usually inevitable to make modifications and re-train the original prediction model. In comparison, the proposed method in this paper is a straightforward yet powerful approach to generalization improvement by establishing a routing function and incorporating a rule-based baseline prediction model. The routing function is trained concurrently with the prediction model and decides on whether to switch to the rule-based model when the learning-based prediction model is unreliable.
@@
II Related Works
II-A Motion Prediction with Uncertainty Estimation
Motion prediction algorithms for autonomous driving have been successful on many datasets, and have been integrated into the autonomy stack [21, 22, 23, 24, 25]. However, it is not rare that prediction failure causes erroneous downstream motion planning for autonomous vehicles [26]. Therefore, it is desired to detect such prediction failure in an efficient yet reliable manner. There have been many efforts to leverage uncertainty estimation to decide whether a prediction is reliable [27]. The prediction uncertainty was estimated by various methods, including ensemble [28, 29, 30], dedicated uncertainty estimation model training [27], rule-based estimation [31], and data augmentation [32]. However, training new models specialized in uncertainty estimation and performing evaluations on its accuracy is not a straightforward task, because there is often no ground truth. Ensemble-based uncertainty estimation is costly both during training and inference and may introduce too much variance, reducing the reliability of out-of-distribution detection as we show in the ablation study in Sec. V-F. Our method of training a routing function concurrently with individual learning-based predictors can increase the exposure of the routing function to anomalous trajectory prediction upon the normal training dataset, and therefore the final prediction selected from various predictor experts can have better performance on zero-shot generalization tasks.
@@ II-B Mixture-of-Experts
There are also mixture-of-experts methods that collect a set of experts specializing in different sub-tasks, which are likely to be included in the target domain [8]. These methods will then choose one suitable expert to be activated during inference. In our setting, we do not assume a pre-defined set of sub-tasks in the target domain, and we also observed that deep learning-based predictors tend to have unsatisfying performance on cross-dataset generalization. Therefore, we follow the idea of mixture-of-experts but do not train individual experts for specific sub-tasks. We include both deep learning-based and rule-based experts that can perform general motion prediction tasks. A routing function is trained concurrently with deep learning-based predictors, so it is exposed to more diverse trajectory prediction candidates and hence the difficulty of ranking anomalous predictions is mitigated.
@@ III Problem Formulation
In this paper, we focus on zero-shot learning and evaluate the motion prediction neural network models on samples that were not observed during training in the autonomous driving domain. Specifically, we denote ${\bm{x}}_{i}^{1:T}=\{{\bm{x}}_{i}^{t}|t\in\{1,\dots,T\}\}$ as a single agent trajectory in the $i$-th scene, represented by a series of features ${\bm{x}}_{i}^{t}$ from timestep $1$ to $T$. The agents are constantly interacting with the environment for which the context information can be represented by ${\bm{c}}_{i}^{1:T}=\{{\bm{c}}_{i}^{t}|t\in(1,T)\}$. The $i$-th scene is denoted by ${\bm{s}}_{i}=\{({\bm{x}}_{i}^{t},{\bm{c}}_{i}^{t})|t\in(1,T)\}$. The task of the prediction model is to predict future trajectory distribution $p({\bm{x}}_{i}^{T_{\text{h}}+1:T_{\text{f}}}|{\bm{x}}_{i}^{1:T_{\text{h}}},{\bm{c}}_{i}^{1:T_{\text{h}}})$ for an ego agent given its history features (states) ${\bm{x}}_{i}^{1:T_{\text{h}}}$ and context information ${\bm{c}}_{i}^{1:T_{\text{h}}}$ in the $i-$th scene, where $T_{\text{h}}$ is the history horizon and $T_{\text{f}}$ is the lookahead horizon and $T=T_{\text{h}}+T_{\text{f}}$.
@@
We are particularly interested in inspecting and improving the generalization ability upon deep learning-based prediction model, which is trained on one dataset $\mathcal{D_{T}}=\{{\bm{s}}_{i}|i\in(1,M_{\mathcal{T}})\}$, and evaluated on another dataset $\mathcal{D_{E}}=\{{\bm{s}}_{i}|i\in(1,M_{\mathcal{E}})\}$. Note that in this paper, the training and the evaluation datasets are defined differently than the normal convention of training and validation. They may or may not be generated from the same underlying distribution. We evaluate the trained models in a completely new dataset, e.g., training on WOMD and testing on Argoverse.
@@ IV-C Learned Routing Function
As the deep learning-based prediction model is being trained, we collect all its multi-modal prediction outputs. These outputs and the predictions of the rule-based experts for the same agent in the same scene are paired and both compared against the ground truth trajectory in terms of some metric, e.g., the average displacement error. Therefore, we can have a ground truth of which predicted trajectory is better among the two. These pairs and labels are stored in a new data buffer than the original training dataset and are used to train the learned routing function. As the training of the transformer prediction model goes on, its prediction output goes from sub-optimal to more reasonable than the rule-based expert predictions. Thus, the learned routing function can have access to both cases where transformer prediction is worse or better than the rule-based prediction, and hence we can avoid the issue of mode collapse.
@@ V-A1 Deep Learning Prediction Expert
We adopt the state-of-the-art prediction architecture MotionTransformer (MTR) [2] as the backbone of the deep learning prediction expert. It ingests a vectorized representation, including both history trajectories of the vehicles in the scene and road map polylines, as the input representation [36], where all the vector inputs are centered around the ego agent. The inputs are first preprocessed by a PointNet-like [37] polyline encoder and then fed into the transformer scene context encoder. The scene encoder enforces local attention which emphasizes the focus on local context information by adopting $k-$nearest neighbor to find $k$ closest polylines to the polyline of interest. The scene context encoded by the local scene encoder is then enhanced by a dense future prediction, containing future interaction information. A static intention and dynamic searching query pair is generated and input to the scene decoder, along with the enhanced scene context encoding and a query content feature. A prediction head is applied to each decoder layer to generate future trajectories, which are represented by a Gaussian Mixture Model to capture multimodal agent behaviors. Please refer to [2] for more model details.
@@ V-A2 Rule-Based Prediction Expert
As described in Sec. IV-B, we apply a constant velocity model as the rule-based prediction expert. It is also possible to adopt other more complicated rule-based prediction models to incorporate more information such as lane and traffic rules. However, this choice is to demonstrate that even a basic complement to a deep learning model can improve the generalization ability of the whole algorithm.
@@ V-B Baselines
  • MTR: MTR trained on one dataset (WOMD or Argoverse) and zero-shot tested on another.
  • MTR (Oracle): MTR trained and tested on the same dataset, which serves as the performance upper bound.
@@
  • APE (BS): uses an alternate routing function based on the variance of an ensemble of MTR models. Refer to section V-F for more details.
@@ V-C Evalution Metrics
  • minADE: This metric computes the average of the $l2-$displacement between the ground truth trajectory and the closest prediction among six trajectory predictions:
    $\text{minADE}=\min_{k\in\{1,\dots,\kappa\}}\dfrac{1}{T}\sum_{t=1}^{T}||\bar{s}_{G}^{t}-s_{G}^{k,t}||_{2}.$
    where $\kappa$ represents the number of modes predicted by the model. We choose $\kappa=6$ in this paper. We also use this metric to rank the prediction generated by different experts when training the learned routing function.
  • minFDE: This metric computes the $l2-$displacement between the ground truth trajectory and the closest prediction at the last time step:
    $\text{minFDE}=\min_{k\in\{1,\dots,\kappa\}}||\bar{s}_{G}^{T}-s_{G}^{k,T}||_{2}.$
@@
  • mAP: This metric is computed on top of Miss Rate. The predictions that are classified as a miss are labeled as negative, whereas non-miss is positive. Precision and recall are computed based on sorted confidence scores associated with each prediction. The mAP metric is then computed as the interpolated precision values in [39]. This metric offers a holistic assessment of the motion prediction performance.
In addition to the aforementioned common metrics for prediction tasks, we also propose to use Performance Gain Percentage to quantify the improvement upon baseline methods with our proposed method, which is defined as
@@ V-D Implementation Details
The feature input projection layer is set to be a $3-$layer MLP with a hidden dimension of $256$. We stack 6 transformer layers for the scene encoding layer. The embedding feature dimension of these layers is set to be $256$. The motion prediction decoder and output projection head follow the implementation of MTR [2]. The routing function decoder is a stack of $6$ transformer layers with an embedding feature dimension of $256$. The output projection head is a $3-$layer MLP with a hidden dimension of $256$. The routing function decoder and output projection are updated with the scene encoder frozen, after each time the motion prediction decoder is updated. The models are trained by AdamW optimizer on $4$ GPUs (Nvidia RTX $6000$) for $30$ epochs with a batch size of $60$ and a learning rate of $1$e44-4- 4, which is decayed every $2$ epochs by a factor of $0.5$.
@@ V-E Prediction Generalization Performance
Method Validation/Test Train mAP $\uparrow$ minADE $\downarrow$ minFDE $\downarrow$ Miss Rate $\downarrow$ MTR (Oracle) Argoverse Argoverse $0.5135$ $0.4113$ $0.8369$ $0.0567$ MTR Argoverse WOMD $0.1290$ $6.8544$ $12.8778$ $0.6558$ Const-Vel Argoverse - $0.1347$ $3.2680$ $8.0422$ $0.5747$ APE (BS) Argoverse WOMD $0.1319$ $4.1853$ $9.9773$ $0.5912$ Ours (APE) Argoverse WOMD $\mathbf{0.1378}$ $\mathbf{3.0461}$ $\mathbf{7.1423}$ $\mathbf{0.5399}$ MTR (Oracle) WOMD WOMD $0.4477$ $0.7546$ $1.5267$ $0.1529$ MTR WOMD Argoverse $0.0525$ $5.0328$ $9.7935$ $0.7382$ Const-Vel WOMD - $0.0292$ $6.5713$ $16.6447$ $0.8985$ APE (BS) WOMD Argoverse $0.0310$ $6.1452$ $14.9673$ $0.8515$ Ours (APE) WOMD Argoverse $\mathbf{0.0741}$ $\mathbf{4.4099}$ $\mathbf{8.3795}$ $\mathbf{0.6701}$
@@
When generalizing from Argoverse to WOMD, the constant velocity model outperforms MTR in terms of minADE. This shows that it is possible for a rule-based predictor to perform better than a deep learning-based predictor, and hence it is necessary to design strategies to improve the generalization ability of a learning-based predictor, with the proposed APE as one possible solution.

### v2

@@ Abstract
Deep learning-based trajectory prediction models for autonomous driving often struggle with generalization to out-of-distribution (OOD) scenarios, sometimes performing worse than simple rule-based models. To address this limitation, we propose a novel framework, Adaptive Prediction Ensemble (APE), which integrates deep learning and rule-based prediction experts. A learned routing function, trained concurrently with the deep learning model, dynamically selects the most reliable prediction based on the input scenario. Our experiments on large-scale datasets, including Waymo Open Motion Dataset (WOMD) and Argoverse, demonstrate improvement in zero-shot generalization across datasets. We show that our method outperforms individual prediction models and other variants, particularly in long-horizon prediction and scenarios with a high proportion of OOD data. This work highlights the potential of hybrid approaches for robust and generalizable motion prediction in autonomous driving. More details can be found on the project page: https://sites.google.com/view/ape-generalization.
@@ I Introduction
Trajectory prediction is critical for safe and reliable autonomous vehicle systems. Existing prediction algorithms [1, 2, 3, 4] have achieved high accuracy on real-world scenarios, such as real traffic datasets. However, most of these algorithms only work best for in-distribution scenarios. Intuitively, traffic scenarios in different cities of the same country should not possess drastic differences, and human driving skills including their prediction and judgment, are not significantly affected. This is unfortunately not the case for deep learning-based prediction algorithms [5, 6]. If they are applied to out-of-distribution (OOD) scenes in a zero-shot manner, such as predicting vehicle trajectories from a different dataset than the training dataset, the performance will drop dramatically even though the input representation and format are the same. In some cases, a deep learning-based prediction algorithm is not even as good as a simple constant velocity model, as shown in Fig. 1. This is unfortunately a largely under-explored topic. One natural way is to combine the prediction from different sources, which resembles the mixture of experts. As far as we know, we are the first to explore concrete methods to improve OOD generalization to different datasets than training.
@@
Figure 1: Illustration of the motivation of improving prediction algorithm by Adaptive Prediction Ensemble. (a) An example scenario where vanilla rule-based prediction algorithm outperforms deep NN prediction algorithm (MTR [1]). (b) A comparison of the error (minADE) between deep NN and rule-based prediction. The rule-based method outperforms deep NN in a considerable amount of scenarios, which are the ones below the red line.
@@
Mixture of Experts (MoEs) [7] has gained popularity, especially after the great success of Large Language Models. Many prior work showed MoEs can reach faster inference [8, 9] compared to dense models with the same number of parameters, and can also be pre-trained faster [10]. While they focused on MoEs’ advantage over a comparable dense model in size, we are more interested in investigating the generalization ability improvements upon a single expert model. We generally find that deep learning prediction models tend to overfit their training dataset, making zero-shot performance unacceptable. Incorporating a fleet of deep learning prediction experts or adopting similar size large models would not solve the problem, since increasing model capacity would not mitigate the overfitting problem if not worse. Therefore, we propose to employ a rule-based prediction expert as an anomaly-handling strategy for deep learning prediction experts, in light of the insight that rule-based prediction could be more reliable in long-tail cases of deep learning prediction experts.
@@
There are other existing methods for domain generalization that usually handle the problem by data manipulation [11], representation learning [12, 13], or specially designed learning strategy [14, 15, 16] or inference workflow [17]. If one aims to improve the generalization upon an existing prediction model with prior methods, it is usually inevitable to make modifications and re-train the original prediction model. In comparison, the proposed method in this paper is a straightforward yet powerful approach to generalization improvement by establishing a routing function and incorporating a rule-based baseline prediction model. The routing function is trained concurrently with the prediction model and decides on whether to switch to the rule-based model when the learning-based prediction model is unreliable.
@@
II Related Work
II-A Improving Out-of-Distribution Motion Prediction
Motion prediction algorithms for autonomous driving have been successful on many datasets, and have been integrated into the autonomy stack [20, 21, 22, 23, 24]. However, OOD performance is known to drop for machine learning algorithms in general [25, 26], and transformer architectures in specific [27, 28], warranting the need for fine-tuning OOD [29, 30]. This is problematic since deployed models are expected to work everywhere, and it is not rare that prediction failure causes erroneous downstream motion planning for autonomous vehicles [31]. Therefore, it is desired to detect such prediction failure in an efficient yet reliable manner. There have been many efforts to leverage uncertainty estimation to decide whether a prediction is reliable based on ensemble [32, 33]. However, ensemble-based uncertainty estimation is costly both during training and inference and may introduce too much variance, reducing the reliability of out-of-distribution detection as we show in the ablation study in Sec. V-F. Our method of training a routing function concurrently with individual learning-based predictors can increase the exposure of the routing function to anomalous trajectory prediction upon the normal training dataset, and therefore the final prediction selected from various predictor experts can have better performance on zero-shot generalization tasks.
@@ II-B Mixture-of-Experts
There are also mixture-of-experts methods that collect a set of experts specializing in different sub-tasks, which are likely to be included in the target domain [7]. These methods will then choose one suitable expert to be activated during inference. In our setting, we do not assume a pre-defined set of sub-tasks in the target domain, and we also observed that deep learning-based predictors tend to have unsatisfying performance on cross-dataset generalization. Therefore, we follow the idea of MoEs but do not train individual experts for specific sub-tasks. We include both deep learning-based and rule-based experts that can perform general motion prediction tasks. A routing function is trained concurrently with deep learning-based predictors, so it is exposed to more diverse trajectory prediction candidates and hence the difficulty of ranking anomalous predictions is mitigated.
@@ III Problem Formulation
In this paper, we focus on zero-shot learning and evaluate the motion prediction neural network models on samples that were not observed during training in the autonomous driving domain. Specifically, we denote ${\bm{x}}_{i}^{1:T}=\{{\bm{x}}_{i}^{t}|t\in\{1,\dots,T\}\}$ as a single agent trajectory in the $i$-th scene, represented by a series of features ${\bm{x}}_{i}^{t}$ from timestep $1$ to $T$. The agents are constantly interacting with the environment for which the context information can be represented by ${\bm{c}}_{i}^{1:T}=\{{\bm{c}}_{i}^{t}|t\in(1,T)\}$. The context information includes map polylines and surrounding agent polylines, which are represented by a series of vectors containing coordinates, direction, etc. The $i$-th scene is denoted by ${\bm{s}}_{i}=\{({\bm{x}}_{i}^{t},{\bm{c}}_{i}^{t})|t\in(1,T)\}$. The task of the prediction model is to predict future trajectory distribution $p({\bm{x}}_{i}^{T_{\text{h}}+1:T_{\text{f}}}|{\bm{x}}_{i}^{1:T_{\text{h}}},{\bm{c}}_{i}^{1:T_{\text{h}}})$ for an ego agent given its history features (states) ${\bm{x}}_{i}^{1:T_{\text{h}}}$ and context information ${\bm{c}}_{i}^{1:T_{\text{h}}}$ in the $i-$th scene, where $T_{\text{h}}$ is the history horizon and $T_{\text{f}}$ is the lookahead horizon and $T=T_{\text{h}}+T_{\text{f}}$.
@@
We are interested in inspecting and improving the generalization ability upon deep learning-based prediction model, which is trained on one dataset $\mathcal{D_{T}}=\{{\bm{s}}_{i}|i\in(1,M_{\mathcal{T}})\}$ with $M_{\mathcal{T}}$ scenes, and evaluated on another dataset $\mathcal{D_{E}}=\{{\bm{s}}_{i}|i\in(1,M_{\mathcal{E}})\}$ with $M_{\mathcal{E}}$ scenes. Note that in this paper, the training and the evaluation datasets are defined differently than the normal convention of training and validation. They may or may not be generated from the same underlying distribution. We evaluate the trained models in a completely new dataset, e.g., training on WOMD and testing on Argoverse.
@@ IV-C Learned Routing Function
While training the learning-based prediction model, we collect all its multi-modal prediction outputs. These outputs and the predictions of the rule-based experts for the same agent in the same scene are paired and both compared against the ground truth trajectory in terms of some metric, e.g., the average displacement error. Therefore, we can have a ground truth of which predicted trajectory is better between the two. These pairs and labels are stored in a new data buffer than the original training dataset and are used to train the learned routing function. As the training of the transformer prediction model goes on, its prediction output goes from sub-optimal to more reasonable than the rule-based expert predictions. Thus, the learned routing function can have access to both cases where transformer is worse or better than the rule-based prediction, and hence we can avoid the issue of mode collapse.
@@ V-A1 Deep Learning Prediction Expert
We adopt the state-of-the-art prediction architecture MotionTransformer (MTR) [1] as the backbone of the deep learning prediction expert. It ingests a vectorized representation, including both history trajectories of the vehicles in the scene and road map polylines, as the input representation [37], where all the vector inputs are centered around the ego agent. The inputs are first preprocessed by a PointNet-like [38] polyline encoder and then fed into the transformer scene context encoder. The scene encoder enforces local attention which emphasizes the focus on local context information by adopting $k-$nearest neighbor to find $k$ closest polylines to the polyline of interest. The scene context encoded by the local scene encoder is then enhanced by a dense future prediction, containing future interaction information. A static intention and dynamic searching query pair is generated and input to the scene decoder, along with the enhanced scene context encoding and a query content feature. A prediction head is applied to each decoder layer to generate future trajectories, which are represented by a Gaussian Mixture Model to capture multimodal agent behaviors. Please refer to [1] for more model details. To demonstrate our proposed APE is model agnostic, we also perform experiments based on Wayformer [4], which is another leading motion prediction model based on transformer.
@@ V-A2 Rule-Based Prediction Expert
As described in Sec. IV-B, we apply a constant velocity model as the rule-based expert. It is also possible to adopt other complicated rule-based models to incorporate more information such as lane and traffic rules. However, this choice is to demonstrate that even a basic complement to a deep learning model can improve the generalization ability of the whole algorithm.
@@ V-B Baselines
  • MTR & Wayformer: MTR/Wayformer trained on one dataset (WOMD or Argoverse) and zero-shot tested on another. We follow the Wayformer implementation in UniTraj [40], as the original code is not public released.
  • MTR (in-dist): MTR trained and tested on the same dataset, which serves as the performance upper bound.
@@
  • APE (MTR-bs): uses an alternate routing function based on the variance of an ensemble of MTR models. Refer to section V-F for more details.
@@ V-C Evalution Metrics
  • minADE / minFDE: This metric computes the average or at the last time step of the $l2-$displacement between the ground truth trajectory and the closest prediction among six trajectory predictions [1]. We also use this metric to rank the prediction generated by different experts when training the learned routing function.
@@
  • mAP: This metric is computed on top of Miss Rate, as the interpolated precision values in [41]. It offers a holistic assessment of the motion prediction performance.
We also propose Performance Gain Percentage to quantify the improvement upon baseline methods, which is defined as
@@ V-D Implementation Details
The feature input projection layer is set to be a $3-$layer MLP with a hidden dimension of $256$. We stack 6 transformer layers for the scene encoding layer. The embedding feature dimension of these layers is set to be $256$. The motion prediction decoder and output projection head follow the implementation of MTR [1]. The routing function are $3$ dense layers with an embedding feature dimension of $256$ with multiple projection heads for context information, which are updated with the scene encoder frozen, after each time the motion prediction decoder is updated. The models are trained by AdamW optimizer on $4$ GPUs (Nvidia RTX $6000$) for $30$ epochs with a batch size of $60$ and a learning rate of $1$e44-4- 4, which is decayed every $2$ epochs by a factor of $0.5$.
@@ V-E Prediction Generalization Performance
Method Validation/Test Train mAP $\uparrow$ minADE $\downarrow$ minFDE $\downarrow$ Miss Rate $\downarrow$ MTR (in-dist) Argoverse Argoverse $0.5135$ $0.4113$ $0.8369$ $0.0567$ MTR Argoverse WOMD $0.1290$ $6.8544$ $12.8778$ $0.6558$ Wayformer Argoverse WOMD $0.1275$ $6.8813$ $12.9071$ $0.6648$ Const-Vel Argoverse - $0.1347$ $3.2680$ $8.0422$ $0.5747$ APE (MTR-bs) Argoverse WOMD $0.1319$ $4.1853$ $9.9773$ $0.5912$ APE (MTR) Argoverse WOMD $\mathbf{0.1378}$ $\mathbf{3.0461}$ $\mathbf{7.1423}$ $\mathbf{0.5399}$ APE (Wayformer) Argoverse WOMD $0.1362$ $3.1013$ $7.1968$ $0.5417$ MTR (in-dist) WOMD WOMD $0.4477$ $0.7546$ $1.5267$ $0.1529$ MTR WOMD Argoverse $0.0525$ $5.0328$ $9.7935$ $0.7382$ Wayformer WOMD Argoverse $0.0541$ $5.0266$ $9.7419$ $0.7343$ Const-Vel WOMD - $0.0292$ $6.5713$ $16.6447$ $0.8985$ APE (MTR-bs) WOMD Argoverse $0.0310$ $6.1452$ $14.9673$ $0.8515$ APE (MTR) WOMD Argoverse $\mathbf{0.0741}$ $\mathbf{4.4099}$ $8.3795$ $\mathbf{0.6701}$ APE (Wayformer) WOMD Argoverse $0.0735$ $4.4135$ $\mathbf{8.3706}$ $0.6859$
@@
When generalizing from Argoverse to WOMD, the constant velocity model outperforms MTR in terms of minADE. This shows that it is possible for a rule-based predictor to perform better than a deep learning-based predictor, and hence it is necessary to design strategies to improve the generalization ability of a learning-based predictor, with the proposed APE as one possible solution. We observe the same trend for Wayformer and its APE version. This confirms that APE is model agnostic, which provides generalization benefits to different deep learning based prediction models.
However, there is still potential for improvements compared to the in-distribution performance. The reason is that the learned routing function does not modify individual prediction candidates but chooses one as the final prediction output. Therefore, the upper bound of performance is constrained by the best prediction candidate in a particular scene. To further improve the performance, the bottleneck lies in the individual prediction experts, and we defer to future works to discover more capable individual models that generalize well.